Use counting arguments from the preceding chapter.

My couch potato friend enjoys sitting in front of the TV and grabbing handfuls of 4 chocolates at random from his snack jar. Unbeknownst to him, I have replaced one of the 22 chocolates in his jar with a cashew. (He hates cashews with a passion.)

(a) How many possible outcomes are there the first time he grabs 4 chocolates?
____ outcomes

(b) How many of these include the cashew?
____ outcomes
Answer #1

a) number of outcome to grab 4 chocolates from 21 chocolates =(21C4) =21!/(4!*17!) =5985

b) number of outcome to pick the cashew =N(choose 3 from 21 chocolates and 1 from 1 cashew)

=(21C3)*(1C1) =1330*1 =1330